# Radar-Guided Polynomial Fitting for Metric Depth Estimation

### v2

@@ 3.4 Loss Function
$\mathcal{L}=\lambda_{L_{1}}\|\hat{d}-d_{\text{gt}}\|_{1}+\lambda_{L_{2}}\|\hat{d}-d_{\text{gt}}\|_{2}^{2}+\lambda_{\text{mono}}\left\|\mathbf{1}_{H\times W}-\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}\right\|_{1},$ (6)
where $d_{\text{gt}}\in\mathbb{R}_{+}^{H\times W}$ is the ground truth metric depth map, $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ is our predicted metric depth, and $z\in\mathbb{R}_{+}^{H\times W}$ is the scaleless depth map predicted by the MDE model.
@@
This regularization enforces that our polynomial fitting function remains approximately monotonically increasing, akin to isotonic regression (Barlow et al., 1972). Generally, within an object or local region, a pixel with a higher initial scaleless depth value should not be assigned a lower final metric depth value compared to a pixel with a lower initial scaleless depth. The preservation of ordinal relationships is an inductive bias grounded in the assumption that the MDE model provides a reasonably accurate estimation of intra-object relative depth.
Polynomial fitting introduces a large function space, with expressiveness growing with degree $N$, which can significantly and detrimentally disrupt local monotonic depth ordering if unconstrained. Our regularization term addresses this by preserving local depth ordinality while still allowing corrections of cross-object misalignments. This constraint prevents overfitting of nonlinear transformations to noisy radar data, spurious correlations, and outliers, thereby avoiding the oscillatory behavior that is a known challenge with fitting higher-degree polynomials (Bishop, 2006).
Figure 4: Qualitative results on nuScenes. GET-UP and RadarCam-Depth (RC-D) fail to reconstruct entire regions, yielding objects with large depth errors. Raw MDE yields reasonable relative reconstructions but suffers from incorrect global scale and cross-object misalignments. POLAR leverages polynomial fitting to recover a global scale and correct these misalignments. See Fig. 5 for colorbars.
@@ 4 Experiments
Baseline Methods. We consider five recent radar-camera depth estimation baselines: RadarNet (Singh et al., 2023), SparseBeatsDense (Li et al., 2024b), GET-UP (Sun et al., 2025), RadarCam-Depth (Li et al., 2024a), and TacoDepth (Wang et al., 2025). Notably, both RadarCam-Depth and TacoDepth also take scaleless MDE predictions as input, together with radar points, to predict metric depth.
Datasets. We evaluate all methods on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft (VoD) datasets (see Sec. C for more details) using the MAE and RMSE metrics with maximum evaluation distances of 50, 70, and 80 meters, following established conventions in the literature.
@@ 4.1 Main Results
While RadarCam-Depth also utilizes an MDE backbone, its complex processing of raw MDE predictions can be seen to distort initially correct depth structure. As shown in the top-left highlighted region of Image C, the tree trunk and attached fern are incorrectly placed at different depths despite the initial MDE predictions correctly positioning them at similar depths. Furthermore, overfitting to specific regions within a scene—particularly those with denser ground truth—may reduce overall error relative to the MDE model but can lead to unintended geometric artifacts. This can result in the omission of entire structures in the predicted depth map, such as the highlighted building on the left of Image A and the construction crane arms in Image B. GET-UP also struggles to accurately infer depth structure, resulting in the omission of objects across all three samples: the highlighted building in Image A, the crane arms in Image B, and foreground tree branches in Image C.
As the polynomial degree—selected as a hyperparameter—increases, the maximum possible number of inflection points also grows, giving POLAR greater expressive power to correct misalignments between the many scene elements in Fig. 3, including cars, trucks, trees, buildings, and traffic lights. This is reflected in the successively improved error maps. See Sec. A for quantitative evaluations.
@@ 4.2 Computational Efficiency
Table 2: Training times (minutes per epoch) on nuScenes using an NVIDIA A6000 GPU.
Method Train Time / Epoch
Lin (Lin et al., 2020) 89.25
RadarNet (Singh et al., 2023) 101.50
SparseBeatsDense (Li et al., 2024b) 63.96
GET-UP (Sun et al., 2025) 249.57
RadarCam-Depth (Li et al., 2024a) 86.38
POLAR (Ours) 33.16
Method Inference Time GFLOPs
Lin (Lin et al., 2020) 129.96 550.43
SparseBeatsDense (Li et al., 2024b) 97.47 532.74
GET-UP (Sun et al., 2025) 445.45 630.99
RadarCam-Depth (Li et al., 2024a) 315.64 619.02
TacoDepth (Wang et al., 2025) 29.30 139.87
POLAR (Ours) 24.81 89.70
@@
5 Discussion
Limitations. POLAR requires tuning the polynomial degree as a hyperparameter (Sec. A), which may vary across datasets. Additionally, while our novel first-derivative regularization mitigates oscillatory behavior from high-degree polynomials, further constraints could enhance stability.
Appendix
Appendix A Comparative Studies
Table 4: MDE backbone comparative studies on nuScenes.
nuScenes
Method MAE RMSE
DPT (Ranftl et al., 2021) 5188.2 6884.5
Depth Anything (Yang et al., 2024) 2404.9 4851.1
Depth Pro (Bochkovskii et al., 2024) 3835.0 6600.3
UniDepth (Piccinelli et al., 2025) 2129.8 4887.7
RadarCam-Depth w/ DPT 1689.7 3948.0
RadarCam-Depth w/ Depth Anything 1953.6 5107.8
RadarCam-Depth w/ Depth Pro 3417.1 6462.0
RadarCam-Depth w/ UniDepth 1872.0 4321.2
POLAR w/ DPT 1525.6 3745.0
POLAR w/ Depth Anything 1515.1 3719.4
POLAR w/ Depth Pro 1627.7 4143.9
POLAR w/ UniDepth 1407.8 3193.5
MDE Models. We evaluate POLAR using four monocular depth estimation (MDE) backbones: DPT (Ranftl et al., 2021), Depth Anything (Yang et al., 2024), Depth Pro (Bochkovskii et al., 2024), and UniDepth (Piccinelli et al., 2025). DPT and Depth Anything infer scaleless inverse depth, and while Depth Pro and UniDepth output metric depth, their raw predictions exhibit significant deviation from metric ground truth (see raw MDE performance in Tabs. 41010), and thus we process them the same way as scaleless depth.
Table 5: Sensitivity study of polynomial degree, selected as a hyperparameter.
@@ Appendix A Comparative Studies
Table 6: Ablation studies of architecture and loss components.
@@
cross-modality att. 2238.8 4817.4 2147.9 4926.1
@@
pos. embeddings 1454.1 3488.9 1500.1 3951.8
no ablations 1407.8 3193.5 629.6 1171.3
Polynomial Degree. Tab. 6 quantifies the performance gains achieved with higher-degree polynomials, with degree 8 outperforming lower degrees. At degree 10, performance degrades slightly, potentially due to excessive flexibility resulting in detrimental oscillations.
Figure 5: To quantify absolute improvement in our qualitative results, we provide the colorbars that were used in all examples in Figs. 346, and 7.
Ablations. We evaluate the impact of key components of POLAR in Tab. 6. Replacing learnable prototypes with self-attention on the radar point features degrades performance, demonstrating that prototypes capture relevant patterns within radar point configurations. Removing cross-modality fusion results in the largest performance drop, highlighting the necessity of leveraging correspondences between MDE and radar features. Removing the monotonicity loss term also reduces performance, suggesting that depth ordering regularization is crucial for stable polynomial fitting.
@@ Appendix D Full nuScenes Benchmark
Table 7: Full quantitative results ($m$) on the nuScenes benchmark.
@@
50m NLSPN (Park et al., 2020) 2790.0 5813.4
BPNet (Tang et al., 2024) 2407.0 4438.0
RC-PDA (Long et al., 2021) 2225.0 4156.5
RC-PDA-HG (Long et al., 2021) 2210.0 4234.0
BTS (Lee et al., 2019) 1937.0 3885.0
DORN (Lo & Vandewalle, 2021) 1926.6 4124.8
RadarNet (Singh et al., 2023) 1727.7 3746.8
CaFNet (Sun et al., 2024) 1674.0 3674.0
Lin (Lin et al., 2020) 1598.2 3790.1
SparseBeatsDense (Li et al., 2024b) 1524.5 3567.3
RadarCam-Depth (Li et al., 2024a) 1286.1 2964.3
GET-UP (Sun et al., 2025) 1241.0 2857.0
TacoDepth (Wang et al., 2025) 1046.8 2487.5
@@
70m NLSPN (Park et al., 2020) 3140.0 6580.6
RC-PDA (Long et al., 2021) 3326.1 6700.6
RC-PDA-HG (Long et al., 2021) 3485.6 7002.9
BTS (Lee et al., 2019) 2346.0 4811.0
DORN (Lo & Vandewalle, 2021) 2170.0 4532.0
RadarNet (Singh et al., 2023) 2073.2 4590.7
CaFNet (Sun et al., 2024) 2010.0 4493.0
Lin (Lin et al., 2020) 1897.8 4558.7
SparseBeatsDense (Li et al., 2024b) 1822.9 4303.6
RadarCam-Depth (Li et al., 2024a) 1587.9 3662.5
GET-UP (Sun et al., 2025) 1541.0 3657.0
TacoDepth (Wang et al., 2025) 1347.1 3152.8
@@
80m NLSPN (Park et al., 2020) 3257.7 6872.4
RC-PDA (Long et al., 2021) 3721.0 7632.0
RC-PDA-HG (Long et al., 2021) 3664.0 7775.0
AdaBins (Bhat et al., 2021) 3541.0 5885.0
P3Depth (Patil et al., 2022) 3130.0 5838.0
LapDepth (Song et al., 2021) 2544.0 5151.0
PnP (Wang et al., 2018) 2496.0 5578.0
BTS (Lee et al., 2019) 2467.0 5125.0
DORN (Lo & Vandewalle, 2021) 2432.0 5304.0
RadarNet (Singh et al., 2023) 2179.3 4898.7
CaFNet (Sun et al., 2024) 2109.0 4765.0
Lin (Lin et al., 2020) 1988.4 4841.1
SparseBeatsDense (Li et al., 2024b) 1927.0 4609.6
RadarCam-Depth (Li et al., 2024a) 1689.7 3948.0
GET-UP (Sun et al., 2025) 1632.0 3932.0
TacoDepth (Wang et al., 2025) 1492.4 3324.8
@@ Appendix E Comparison to Depth Completion
One potential idea for radar-camera depth estimation is to apply lidar-camera depth completion methods designed to densify sparse depth maps using surrounding context. One such method, BPNet (Tang et al., 2024) achieves state-of-the-art performance on the KITTI depth completion benchmark by leveraging bilateral propagation. However, when applied to radar-camera depth estimation, BPNet performs poorly, as shown in Tab. 7. POLAR outperforms BPNet by 57.9% in MAE and 44.2% in RMSE on nuScenes, highlighting the limitations of directly applying lidar depth completion methods to radar data. The key reason for this underperformance lies in the fundamental differences between lidar and radar point clouds. Unlike lidar, radar measurements are orders of magnitude sparser and significantly noisier due to factors such as limited antenna elements, elevation ambiguity (see Fig. 7), and lower range resolution (Singh et al., 2023). Lidar depth completion methods assume a relatively dense and structured input (Xia et al., 2023), leveraging local spatial continuity to propagate depth estimates effectively. In contrast, radar points are too sparse for such methods to infer meaningful local depth relationships, leading to poor depth reconstruction when attempting direct densification.
@@
In addition, we compare against Non-Local Spatial Propagation Network (NLSPN) (Park et al., 2020), which achieves near state-of-the-art performance on the KITTI depth completion benchmark by refining sparse lidar depth with an iterative non-local spatial propagation procedure. NLSPN predicts an initial depth map along with pixel-wise confidences, then refines it by estimating non-local neighbors and their corresponding affinities to selectively propagate depth information. Unlike other approaches that rely on fixed local neighbors, NLSPN adaptively determines relevant non-local neighbors, improving depth completion accuracy, especially near depth boundaries. However, again, when applied to radar-camera depth estimation, NLSPN performs poorly, as shown in Tab. 7. POLAR outperforms NLSPN by 59.8% in MAE and 55.4% in RMSE on nuScenes.
@@
Appendix F Evaluation Metrics and Implementation Details
The evaluation metrics used in our study, Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE), are formulated in Tab. 8. Lower values equal better performance for both MAE and RMSE. Unless specified otherwise, all reported values are in millimeters (mm). We train for 60 epochs using a cosine decay learning rate scheduler with learning rate of $5\times 10^{-5}$, and use weighting terms $\lambda_{L_{1}}=1.0,\lambda_{L_{2}}=0.4,\lambda_{mono}=0.25$ for our loss function.
Appendix G Additional Comparisons
Full MDE Comparisons. For ZJU (see Tab. 10), among all configurations, POLAR w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 61.1% and over RadarCam-Depth w/ UniDepth by 54.2%. For VoD (see Tab. 10), POLAR w/ UniDepth achieves the best performance in MAE, improving over raw UniDepth predictions by 42.4% and over RadarCam-Depth w/ UniDepth by 32.7%, while POLAR w/ Depth Anything achieves the best performance in RMSE, improving over RadarCam-Depth w/ Depth Anything by 37.6% and over raw Depth Anything predictions (inverted and median scaled) by 10.4%.
Table 9: MDE backbone comparative studies on ZJU-4DRadarCam (ZJU).
ZJU
Method MAE RMSE
DPT (Ranftl et al., 2021) 1885.3 3326.1
Depth Anything (Yang et al., 2024) 1943.2 3469.3
Depth Pro (Bochkovskii et al., 2024) 1680.2 3144.9
UniDepth (Piccinelli et al., 2025) 1533.0 3188.4
RadarCam-Depth w/ DPT 1183.5 3229.0
RadarCam-Depth w/ Depth Anything 1724.4 3661.3
RadarCam-Depth w/ Depth Pro 1490.6 3429.5
RadarCam-Depth w/ UniDepth 1152.5 3168.6
POLAR w/ DPT 707.1 1216.9
POLAR w/ Depth Anything 657.2 1225.4
POLAR w/ Depth Pro 640.3 1174.8
POLAR w/ UniDepth 629.6 1171.3
Table 10: MDE backbone comparative studies on View-of-Delft.

### v3

@@ 3.4 Loss Function
$\mathcal{L}=\lambda_{1}\|\hat{d}-d\|_{1}+\lambda_{2}\|\hat{d}-d\|_{2}^{2}+\lambda_{\text{m}}\left\|\mathbf{1}_{H\times W}-\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}\right\|_{1},$ (6)
where $d\in\mathbb{R}_{+}^{H\times W}$ is the ground truth metric depth map, $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ is our predicted metric depth, and $z\in\mathbb{R}_{+}^{H\times W}$ is the scaleless depth map predicted by the MDE model.
@@
This regularization enforces that our polynomial fitting function remains approximately monotonically increasing, akin to isotonic regression [barlow1972isotonic]. Generally, within an object or local region, a pixel with a higher initial scaleless depth value should not be assigned a lower final metric depth value compared to a pixel with a lower initial scaleless depth. The preservation of ordinal relationships is an inductive bias grounded in the assumption that the MDE model provides a reasonably accurate estimation of intra-object relative depth.
Polynomial fitting introduces a large function space, with expressiveness growing with degree $N$, which can significantly and detrimentally disrupt local monotonic depth ordering if unconstrained. Our regularization term addresses this by preserving local depth ordinality while still allowing corrections of cross-object misalignments. This constraint prevents overfitting of nonlinear transformations to noisy radar data, spurious correlations, and outliers, thereby avoiding the oscillatory behavior that is a known challenge with fitting higher-degree polynomials [bishop2006pattern].
@@ 4 Experiments
Baseline Methods. We consider five recent radar-camera depth estimation baselines: RadarNet [singh2023depth], SparseBeatsDense [li2023sparse], GET-UP [sun2024get], RadarCam-Depth [li2024radarcam], and TacoDepth [wang2025tacodepth]. Notably, both RadarCam-Depth and TacoDepth also take scaleless MDE predictions as input, together with radar points, to predict metric depth.
Datasets. We evaluate all methods on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft (VoD) datasets (see Supp. Mat. for more details) using the MAE and RMSE metrics with maximum evaluation distances of 50, 70, and 80 meters, following established conventions in the literature.
@@ 4.1 Main Results
RadarCam-Depth predicts a per-pixel scale map; the large number of degrees of freedom can make it challenging to preserve initially correct MDE depth structure. As shown in the top-left of Image C, the tree trunk and attached fern are incorrectly placed at different depths despite the initial MDE predictions correctly positioning them at similar depths. Furthermore, overfitting to specific regions within a scene—particularly those with denser ground truth—may reduce overall error relative to the MDE model but can lead to unintended geometric artifacts. This can result in the omission of entire structures in the predicted depth map, such as the highlighted building on the left of Image A and the crane arms in Image B. GET-UP also struggles to accurately infer depth structure, resulting in the omission of objects: the highlighted building in Image A, the crane arms in Image B, and foreground tree branches in Image C.
As the degree of the polynomial increases, the maximum possible number of inflection points also grows, giving POLAR greater expressive power to correct misalignments between the many elements of the scene in Fig. 4, including cars, trucks, trees, buildings, and traffic lights. This is reflected in the successively improved error maps. See Sec. 5 for quantitative evaluations.
Method Train Time / Epoch
Lin [lin2020depth] 89.25
RadarNet [singh2023depth] 101.50
SparseBeatsDense [li2023sparse] 63.96
GET-UP [sun2024get] 249.57
RadarCam-Depth [li2024radarcam] 86.38
POLAR (Ours) 33.16
Table 2: Training times (minutes per epoch) on nuScenes using a single NVIDIA A6000 GPU.
Method Inference Time GFLOPs
Lin [lin2020depth] 129.96 550.43
SparseBeatsDense [li2023sparse] 97.47 532.74
GET-UP [sun2024get] 445.45 630.99
RadarCam-Depth [li2024radarcam] 315.64 619.02
TacoDepth [wang2025tacodepth] 29.30 139.87
POLAR (Ours) 24.81 89.70
@@ 4.2 Computational Efficiency
5 Comparative Studies
Polynomial Degree. Tab. 4 shows that POLAR with polynomial degree 8 outperforms lower degrees. At degree 10, performance degrades slightly, potentially due to excessive flexibility resulting in detrimental oscillations.
Ablations. As shown in Tab. 5, replacing learnable prototypes with self-attention on the radar point features, or replacing our feature aggregation method (Eq. 2) with cross-attention, both degrade performance, demonstrating that prototypes capture relevant patterns within radar point configurations. Removing the monotonicity loss term also reduces performance, suggesting that depth ordering regularization is crucial for stable polynomial fitting. Finally, learning to directly regress depth from our final fused representation using a decoder head performs substantially worse. We hypothesize that this is because the number of degrees of freedom ($H\times W$ pixels) is far larger than the available constraints (sparse radar points) to solve for depth at every pixel.
@@ 5 Comparative Studies
Table 4: Sensitivity study of polynomial degree hyperparameter.
@@
feature aggregation 1454.1 3488.9 1543.1 4070.6
direct decoding 1968.0 4155.7 1855.9 4608.0
@@
no ablations 1407.8 3193.5 1500.1 3951.8
Table 5: Ablation studies of architecture and loss components.
6 Discussion
Limitations. POLAR requires tuning the polynomial degree as a hyperparameter (Sec. 5), which may vary across datasets. Additionally, while our novel first-derivative regularization mitigates oscillatory behavior from high-degree polynomials, further constraints could enhance stability.
[…]
Supplementary Material
nuScenes
Method MAE RMSE
DPT [ranftl2021vision] 5188.2 6884.5
UniDepth [piccinelli2025unidepthv2] 2129.8 4887.7
Depth Anything [depth_anything_v2] 2404.9 4851.1
Depth Pro [depth-pro] 3835.0 6600.3
RadarCam-Depth w/ DPT 1689.7 3948.0
RadarCam-Depth w/ UniDepth 1872.0 4321.2
RadarCam-Depth w/ Depth Anything 1953.6 5107.8
RadarCam-Depth w/ Depth Pro 3417.1 6462.0
TacoDepth w/ DPT 1492.4 3324.8
POLAR w/ DPT 1447.2 3304.9
POLAR w/ UniDepth 1407.8 3193.5
POLAR w/ Depth Anything 1515.1 3580.0
POLAR w/ Depth Pro 1461.5 4143.9
Table 6: MDE backbone comparative studies on nuScenes.
Appendix A Comparative Studies
MDE Models. We evaluate POLAR using four monocular depth estimation (MDE) backbones: DPT [ranftl2021vision], Depth Anything [depth_anything_v2], Depth Pro [depth-pro], and UniDepth [piccinelli2025unidepthv2]. DPT and Depth Anything infer scaleless inverse depth, and while Depth Pro and UniDepth output metric depth, their raw predictions exhibit significant deviation from metric ground truth (see raw MDE performance in Tabs. 6711), and thus we process them the same way as scaleless depth.
For ZJU (see Tab. 7), among all configurations, POLAR w/ UniDepth achieves the best performance, improving over raw UniDepth predictions by 61.1% and over RadarCam-Depth w/ UniDepth by 54.2%.
TacoDepth. Although TacoDepth experiments with only DPT as its MDE backbone, our experiments show that a different backbone provides further performance gains. Nevertheless, POLAR outperforms TacoDepth by over 3% in MAE when both methods use the same DPT backbone. A similar trend holds on ZJU: even when both methods use the same DPT backbone, POLAR outperforms TacoDepth by 31.5% in MAE, further confirming that POLAR’s state-of-the-art performance is not due to backbone choice.
ZJU
Method MAE RMSE
DPT [ranftl2021vision] 1885.3 3326.1
UniDepth [piccinelli2025unidepthv2] 1533.0 3188.4
Depth Anything [depth_anything_v2] 1943.2 3469.3
Depth Pro [depth-pro] 1680.2 3144.9
RadarCam-Depth w/ DPT 1183.5 3229.0
RadarCam-Depth w/ UniDepth 1152.5 3168.6
RadarCam-Depth w/ Depth Anything 1724.4 3661.3
RadarCam-Depth w/ Depth Pro 1490.6 3429.5
TacoDepth w/ DPT 1032.5 2850.3
POLAR w/ DPT 707.1 1216.9
POLAR w/ UniDepth 629.6 1171.3
POLAR w/ Depth Anything 657.2 1225.4
POLAR w/ Depth Pro 640.3 1174.8
Table 7: MDE backbone comparative studies on ZJU.
Figure 5: To quantify absolute improvement in our qualitative results, we provide the colorbars that were used in all examples in Figs. 6 and 7, as well as Figs. 3 and 4 in the main paper.
@@ Appendix B More on Computational Efficiency
Cost breakdown. Following the reporting protocol of TacoDepth [wang2025tacodepth], we measure inference time as the latency added on top of the MDE backbone. Tab. 8 shows that, even when considering the combined cost, our method remains state-of-the-art in both inference time and GFLOPs.
@@ Appendix C Dataset Details
nuScenes
Method Inference Time GFLOPs
Depth Anything [depth_anything_v2] 28.70 201.03
+ Additional over Depth Anything 315.64 619.02
RadarCam-Depth w/ Depth Anything 344.34 820.05
+ Additional over Depth Anything 29.30 139.87
TacoDepth w/ Depth Anything 58.00 340.90
+ Additional over Depth Anything 24.81 89.70
POLAR w/ Depth Anything 53.51 290.73
Table 8: Computational cost breakdown for methods that leverage MDE. Inference time is measured in milliseconds (ms).
@@ Appendix D Full nuScenes Benchmark
50m RC-PDA [long2021full] 2225.0 4156.5
RC-PDA-HG [long2021full] 2210.0 4234.0
NLSPN [Park2020NonLocalSP] 2185.7 4411.6
BTS [lee2019big] 1937.0 3885.0
DORN [lo2021depth] 1926.6 4124.8
RadarNet [singh2023depth] 1727.7 3746.8
CaFNet [sun2024cafnet] 1674.0 3674.0
BPNet [tang2024bilateral] 1618.9 3596.1
Lin [lin2020depth] 1598.2 3790.1
SparseBeatsDense [li2023sparse] 1524.5 3567.3
RadarCam-Depth [li2024radarcam] 1286.1 2964.3
GET-UP [sun2024get] 1241.0 2857.0
TacoDepth [wang2025tacodepth] 1046.8 2487.5
@@
70m RC-PDA [long2021full] 3326.1 6700.6
RC-PDA-HG [long2021full] 3485.6 7002.9
BTS [lee2019big] 2346.0 4811.0
NLSPN [Park2020NonLocalSP] 2260.9 4645.0
DORN [lo2021depth] 2170.0 4532.0
RadarNet [singh2023depth] 2073.2 4590.7
CaFNet [sun2024cafnet] 2010.0 4493.0
Lin [lin2020depth] 1897.8 4558.7
SparseBeatsDense [li2023sparse] 1822.9 4303.6
RadarCam-Depth [li2024radarcam] 1587.9 3662.5
GET-UP [sun2024get] 1541.0 3657.0
TacoDepth [wang2025tacodepth] 1347.1 3152.8
@@
80m RC-PDA [long2021full] 3721.0 7632.0
RC-PDA-HG [long2021full] 3664.0 7775.0
AdaBins [bhat2021adabins] 3541.0 5885.0
P3Depth [patil2022p3depth] 3130.0 5838.0
LapDepth [song2021monocular] 2544.0 5151.0
PnP [wang2018plug] 2496.0 5578.0
NLSPN [Park2020NonLocalSP] 2519.3 5033.7
BTS [lee2019big] 2467.0 5125.0
DORN [lo2021depth] 2432.0 5304.0
RadarNet [singh2023depth] 2179.3 4898.7
CaFNet [sun2024cafnet] 2109.0 4765.0
Lin [lin2020depth] 1988.4 4841.1
SparseBeatsDense [li2023sparse] 1927.0 4609.6
RadarCam-Depth [li2024radarcam] 1689.7 3948.0
GET-UP [sun2024get] 1632.0 3932.0
TacoDepth [wang2025tacodepth] 1492.4 3324.8
@@
Table 9: Full quantitative results ($m$) on the nuScenes benchmark.
@@ Appendix E Comparison to Depth Completion
One potential idea for radar-camera depth estimation is to apply lidar-camera depth completion methods designed to densify sparse depth maps using surrounding context. One such method, BPNet [tang2024bilateral] achieves state-of-the-art performance on the KITTI depth completion benchmark by leveraging bilateral propagation. However, when applied to radar-camera depth estimation, BPNet performs poorly, as shown in Tab. 9. POLAR outperforms BPNet by 37.3% in MAE and 31.2% in RMSE on nuScenes, highlighting the limitations of directly applying lidar depth completion methods to radar data. The key reason for this underperformance lies in the fundamental differences between lidar and radar point clouds. Unlike lidar, radar measurements are orders of magnitude sparser and significantly noisier due to factors such as limited antenna elements, elevation ambiguity (see Fig. 7), and lower range resolution [singh2023depth]. Lidar depth completion methods assume a relatively dense and structured input [xia2023quadric], leveraging local spatial continuity to propagate depth estimates effectively. In contrast, radar points are too sparse for such methods to infer meaningful local depth relationships, leading to poor depth reconstruction when attempting direct densification.
@@
In addition, we compare against Non-Local Spatial Propagation Network (NLSPN) [Park2020NonLocalSP], which achieves near state-of-the-art performance on the KITTI depth completion benchmark by refining sparse lidar depth with an iterative non-local spatial propagation procedure. NLSPN predicts an initial depth map along with pixel-wise confidences, then refines it by estimating non-local neighbors and their corresponding affinities to selectively propagate depth information. Unlike other approaches that rely on fixed local neighbors, NLSPN adaptively determines relevant non-local neighbors, improving depth completion accuracy, especially near depth boundaries. However, again, when applied to radar-camera depth estimation, NLSPN performs poorly, as shown in Tab. 9. POLAR outperforms NLSPN by 44.1% in MAE and 36.6% in RMSE on nuScenes.
@@
Appendix F Evaluation Metrics and Implementation
Details
The evaluation metrics used in our study, Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE), are formulated in Tab. 10. Lower values equal better performance for both MAE and RMSE. Unless specified otherwise, all reported values are in millimeters (mm). We train for 60 epochs using a cosine decay learning rate scheduler with learning rate of $5\times 10^{-5}$, and use weighting terms $\lambda_{1}=1.0,\lambda_{2}=0.4,\lambda_{m}=0.25$ for our loss function.